OpenStudy (anonymous):

Completing the square 16d^2= 25

OpenStudy (anonymous):

I dont think thats right, because from that info I need to find the vertex

OpenStudy (anonymous):

(4d)^2=25 4d=+/-sqrt(25) d=5/4 and d=-(5/4) I think this should be it?

OpenStudy (anonymous):

It should be like this: \[16d^2 - 25 = 0\] \[(4d)^2 - 5^2 = 0\] \[(4d + 5)(4d - 5) = 0\] \[d = \frac{5}{4} \quad Or \quad d= \frac{-5}{4}\]

OpenStudy (anonymous):

I accidentally bracketed the latter d. My bad.